OpenStudy (anonymous):

Find the side length of the cube with the given volume. V=3 over 125

OpenStudy (anonymous):

fan and medal

OpenStudy (trojanpoem):

The cube volume is given from the following equation: \[V = l^3\] The question mentions that the cube volume is 3/125 so 3/125 = l^3 You can easily solve for l by getting the cubic root.

OpenStudy (anonymous):

so1?

OpenStudy (trojanpoem):

\[l \approx 0.3\]

OpenStudy (anonymous):

I don't get that the answers I can put are 5, 25, 125,and 625

OpenStudy (trojanpoem):

Are you sure that the volume is \[\frac{ 3 }{ 125 }\] (3 over 125} ?